two numbers are in the ratio 4:3. Their sum is 70. Find the numbers.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Their sum is 7
70/7 is 10
it is 10 times each of these
the numbers are 40 and 30, 40:30=4:3

Answer by ikleyn(52780)

.
The fact that the two numbers are in the ratio 4:3 means that the numbers are  4x and 3x, where x is their common measure, 
currently unknown to us.


But we are given that the sum of the numbers is 70, and it gives us an equation


    4x + 3x = 70.


It means that 7x = 70,  which implies  x= 70/7 = 10.


Thus the numbers are 4x = 4*10 = 40  and  3x = 3*10 = 30.     ANSWER

Answered, solved, explained and completed.